# Phase transition near the Big Bang in the lattice theory of gravity and some cosmological considerations

### v6

@@ Abstract
Lattice regularization of the theory of gravity provides a new possibility for study of the Big Bang problem. Moreover, the very possibility of studying the state of the Universe at the highest temperatures appears. In the 4D lattice theories of gravity, the existence of a high-temperature phase is proved, which is characterized by the folding of space into a point and zeroing of the energy-momentum tensor. This fact leads to the hypothesis that there are at least two phases in the 4D lattice theory, and the main properties of the high-temperature phase are similar to those described above. The solution of the Einstein equation in the low temperature phase and the long-wave limit within the framework of the Friedmann paradigm with a finite bare cosmological constant is given. This solution is mathematically correct, since all local physical quantities (energy density including vacuum energy, etc.) on the lattice are finite. As a result, a solution is obtained that demonstrates an exponential growth of the cosmological scale factor $a(t)$ in the initial period of evolution (inflation phase of the Universe) and then passes into a power law ($a(t)\propto\sqrt{t}$). This solution also demonstrates that the energy-momentum tensor tends to zero when approaching the Big Bang point.
@@ I Introduction
As is known, in the Minimum Standard Model (MSM) (considered within the framework of gravity) the number of fermion degrees of freedom is a multiple of the number of bosonic degrees of freedom. Therefore, we assume that some properties of the MSM near the point of the Big Bang can be described using the discrete gravity model being studied. Conversely, the following question is of interest to us: does the continuation of solutions to Einstein’s continuum equations in the vicinity of the Big Bang point contain some information about the properties of the high-temperature phase in lattice theory? To answer this question, we solve Einstein’s equations in the spirit of Friedmann’s model, but with a bare cosmological constant of positive sign. The positive cosmological constant compensates for the enormous energy density of negative-sign quantum zero-point oscillations of the vacuum. Here we make the following remarks: 1) The energy density of zero-point vacuum oscillations is negative, since fermion degrees of freedom predominate in the MSM. 2) We consider the continuum model of gravity with matter fields included as the long-wave limit of the lattice theory. Therefore, all quantities (energy density $\varepsilon$, bare cosmological constant $\Lambda_{0}$, etc.) are finite. In this case, Einstein’s equations and their solutions are mathematically correct. The solution found in this work shows a decrease in the energy density modulus $|\varepsilon|$ to zero when approaching the Big Bang point. Of course, in the vicinity of the Big Bang point, neither the equations nor the solution are correct due to the increase in the value of quantum fluctuations (see Section IV). However, the tendency towards a decrease in $|\varepsilon|$ within the framework of the reliability of the solution indicates the consistency of the lattice model and its long-wave limit. Let us emphasize that the decrease in $|\varepsilon|$ occurs in parallel with an increase in the density of matter.
@@
Section IV considers the solution of the Einstein equation within the Friedmann paradigm in the long-wave limit. The validity of the classical approach (without taking into account quantum fluctuations) at some distance from the Big Bang point is substantiated.
@@ II.1 Lattice theory definition
In this and next Sections, consideration is carried out for the Euclidean signature.
Denote by $\gamma^{a}$ $4\times 4$ Hermitian Dirac matrices, so that
@@
Consider the orientable abstract 4-dimensional simplicial complex $\mathop{{\mathfrak{K}}}\nolimits$. Suppose that any of its 4-simplexes belongs to such a finite (or infinite) sub-complex ${\mathop{{\mathfrak{K}}}\nolimits}^{\prime}\in\mathop{{\mathfrak{K}}}\nolimits$ which has a geometric realization in $\mathbb{R}^{4}$ topologically equivalent to a disk without cavities. The vertices are designated as $a_{{\cal V}}$, the indices ${{\cal V}}=1,2,\dots,\,{\mathop{{\mathfrak{N}}}\nolimits}^{(0)}\rightarrow\infty$ and ${{\cal W}}$ enumerate the vertices and 4-simplices, correspondingly. It is necessary to use the local enumeration of the vertices $a_{{\cal V}}$ attached to a given 4-simplex: the all five vertices of a 4-simplex with index ${{\cal W}}$ are enumerated as $a_{{{\cal V}}_{({{\cal W}})i}}$, $i=1,2,3,4,5$. The later notations with extra low index $({{\cal W}})$ indicate that the corresponding quantities belong to the 4-simplex with index ${{\cal W}}$. Of course, these same quantities also belong to another 4-simplex with index ${{\cal W}}^{\prime}$, and 4-simplexes with indices ${{\cal W}}$ and ${{\cal W}}^{\prime}$ must be adjacent. The Levi-Civita symbol with in pairs different indexes $\varepsilon_{{{\cal V}}_{({{\cal W}})1}{{\cal V}}_{({{\cal W}})2}{{\cal V}}_{({{\cal W}})3}{{\cal V}}_{({{\cal W}})4}{{\cal V}}_{({{\cal W}})5}}=\pm 1$ depending on whether the order of vertices $s^{4}_{{\cal W}}=a_{{{\cal V}}_{({{\cal W}})1}}a_{{{\cal V}}_{({{\cal W}})2}}a_{{{\cal V}}_{({{\cal W}})3}}a_{{{\cal V}}_{({{\cal W}})4}}a_{{{\cal V}}_{({{\cal W}})5}}$ defines the positive or negative orientation of 4-simplex $s^{4}_{{\cal W}}$. An element of the compact group $\mathop{\rm Spin}\nolimits(4)$ and an element of the Clifford algebra
$\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}=\Omega^{-1}_{{{\cal V}}_{2}{{\cal V}}_{1}}=\mathop{\rm exp}\nolimits\left(\omega_{{{\cal V}}_{1}{{\cal V}}_{2}}\right)=\mathop{\rm exp}\nolimits\left(\frac{1}{2}\sigma^{ab}\omega^{ab}_{{{\cal V}}_{1}{{\cal V}}_{2}}\right)\in\mathop{\rm Spin}\nolimits(4),\quad\sigma^{ab}\equiv\frac{1}{4}[\gamma^{a}\gamma^{b}]$
@@
$\mathop{{\mathfrak{A}}}\nolimits_{g}=-\frac{1}{5!\cdot 2\cdot l_{P}^{2}}\sum_{{\cal W}}\sum_{\sigma}\varepsilon_{\sigma({{\cal V}}_{({{\cal W}})1})\sigma({{\cal V}}_{({{\cal W}})2})\sigma({{\cal V}}_{({{\cal W}})3})\sigma({{\cal V}}_{({{\cal W}})4})\sigma({{\cal V}}_{({{\cal W}})5})}$
@@
$\mathop{{\mathfrak{A}}}\nolimits_{\Lambda_{0}}=-\frac{1}{5!\cdot 12}\cdot\frac{\Lambda_{0}}{l_{P}^{2}}\varepsilon_{abcd}\sum_{{\cal W}}\sum_{\sigma}\varepsilon_{\sigma({{\cal V}}_{({{\cal W}})1})\sigma({{\cal V}}_{({{\cal W}})2})\sigma({{\cal V}}_{({{\cal W}})3})\sigma({{\cal V}}_{({{\cal W}})4})\sigma({{\cal V}}_{({{\cal W}})5})}$
@@ II.2 The long-wavelength limit
Now let us pass on to the limit of slowly varying fields, when the action (4) reduces to the well known continuous gravity action. This transition have meaning together with the transition to Minkowski signature. As a result the compact gauge group $\mathop{\rm Spin}\nolimits(4)$ transforms into the non-compact group $\mathop{\rm Spin}\nolimits(3,1)$.
In this section, all lattice quantities for the Euclidean signature are primed. For the quantities related to the Minkowski signature, the previous designations are retained. This is done to make the transition from the Euclidean signature to the Minkowski signature transparent.
Firstly let us perform the following deformations of integration contours in integral (10):
@@
Now let us pass on to the formal limit of slowly varying fields, when the action (4) reduces to the well known continuous gravity action. In order to pass to the long-wavelength limit let’s consider a certain 4D sub-complex ${\mathop{{\mathfrak{K}}}\nolimits}^{\prime}\in\mathop{{\mathfrak{K}}}\nolimits$ with the trivial topology of four-dimensional disk. Realize geometrically this sub-complex in $\mathbb{R}^{4}$. Suppose that the geometric realization is the triangulation of a compact part of $\mathbb{R}^{4}$, so that in $\mathbb{R}^{4}$ we have $(\mbox{int}\,s^{4}_{{\cal W}})\cap(\mbox{int}\,s^{4}_{{{\cal W}}^{\prime}})=\emptyset$ for ${{\cal W}}\neq{{\cal W}}^{\prime}$ and the sizes of these simplices are commensurable. Thus each vertex of the sub-complex acquires the coordinates $x^{\mu}$ which are the coordinates of the vertex image in $\mathbb{R}^{4}$:
@@
We stress that the correspondence between the vertices from the considered subset and the coordinates (24) is one-to-one.
Consider a certain simplex $s^{4}_{{\cal W}}$. We denote all five vertices of this 4-simplex as ${{\cal V}}_{i}$, $i=1,2,3,4$ and ${{\cal V}}_{m}\neq{{\cal V}}_{i}$. Note that each of these vertices also belongs to other adjacent 4-simplexes. From the aforesaid it is evident that the four infinitesimal vectors
@@
are linearly independent. The vector system (25) refers to the simplex $s^{4}_{{\cal W}}$ and its vertex ${{\cal V}}_{m}$, while each of the vectors of the system (25) belongs to several adjacent 4-simplexes. The differentials of coordinates (25) correspond to one-dimensional simplices $a_{{{\cal V}}_{m}}a_{{{\cal V}}_{i}}$, so that, if the vertex $a_{{{\cal V}}_{m}}$ has coordinates $x^{\mu}_{{{\cal V}}_{m}}$, then the vertex $a_{{{\cal V}}_{i}}$ has the coordinates $x^{\mu}_{{{\cal V}}_{m}}+\mathop{\rm d}\nolimits x^{\mu}_{\left({{\cal V}}_{m}{{\cal V}}_{i}\right)}$.
In the continuous limit, the holonomy group elements $\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}$ are close to the identity element, so that the quantities $\omega^{ab}_{{{\cal V}}_{1}{{\cal V}}_{2}}$ tend to zero being of the order of $O(\mathop{\rm d}\nolimits x^{\mu})$ and they slowly change when moving along the lattice. Thus one can consider the following system of equation for $\omega_{\mu}\left({\cal W},{{\cal V}}_{m},{{\cal V}}_{i}\right)$:
[…]
In this system of linear equation, the indices ${{\cal W}}$ and $m$ are fixed, the summation is carried out over the index $\mu$. Since the vectors (25) are linearly independent, the quantities $\omega_{\mu}\left({\cal W},{{\cal V}}_{m},{{\cal V}}_{i}\right)$ are defined uniquely.
The problem is the definition of the connection 1-form field $\omega_{\mu}\big{(}(x_{{{\cal V}}_{m}}+x_{{{\cal V}}_{i}})/2\big{)}$ satisfying systems of equations (26) for all 1-simplexes $a_{{{\cal V}}_{m}}a_{{{\cal V}}_{i}}$ independently of 4-simplexes and the vertices marked in them. The inverse problem is solved obviously: if a 1-form $\omega_{\mu}\big{(}(x_{{{\cal V}}_{m}}+x_{{{\cal V}}_{i}})/2\big{)}$ is given, then the set of values $\omega_{{{\cal V}}_{m}{{\cal V}}_{i}}$ is determined according to (26).
Note that due to Eqs. (2), (25) and (26) we have
$\omega_{\mu}\left({\cal W},{{\cal V}}_{m},{{\cal V}}_{i}\right)\,\mathop{\rm d}\nolimits x^{\mu}_{{{\cal V}}_{m}{{\cal V}}_{i}}=\omega_{\mu}\left({\cal W},{{\cal V}}_{i},{{\cal V}}_{m}\right)\,\mathop{\rm d}\nolimits x^{\mu}_{{{\cal V}}_{m}{{\cal V}}_{i}}.$ (27)
Suppose that a one-dimensional simplex $a_{{{\cal V}}_{m}}a_{{{\cal V}}_{i}}$ belongs to four-dimensional simplices with indices ${{\cal W}}_{1},\,{{\cal W}}_{2},\,\ldots\,,\,{{\cal W}}_{r}$. Introduce the quantity
$\omega_{\mu}\left(\frac{1}{2}\,(x_{{{\cal V}}_{m}}+x_{{{\cal V}}_{i}})\,\right)\equiv\frac{1}{2r}\sum_{s=1}^{r}\bigg{\{}\omega_{\mu}\left({{\cal W}}_{s},{{\cal V}}_{m},{{\cal V}}_{i}\right)+\omega_{\mu}\left({{\cal W}}_{s},{{\cal V}}_{i},{{\cal V}}_{m}\right)\bigg{\}}\,,$ (28)
which is assumed to be related to the midpoint of the segment $[x^{\mu}_{{{\cal V}}_{m}},\,x^{\mu}_{{{\cal V}}_{i}}\,]$. According to the definition, we have the following chain of identities:
$\omega_{{{\cal V}}_{({{\cal W}}_{1})m}{{\cal V}}_{({{\cal W}}_{1})i}}\equiv\ldots\equiv\omega_{{{\cal V}}_{({{\cal W}}_{r})m}{{\cal V}}_{({{\cal W}}_{1})r}}.$ (29)
It follows from (25)–(29) that
$\omega_{\mu}\left(\frac{1}{2}\,(x_{{{\cal V}}_{m}}+x_{{{\cal V}}_{i}})\,\right)\,\mathop{\rm d}\nolimits x^{\mu}_{{{\cal V}}_{m}{{\cal V}}_{i}}=\omega_{{{\cal V}}_{m}{{\cal V}}_{i}}\,.$ (30)
The value of the field element $\omega_{\mu}$ in (30) is uniquely defined by the midpoint of corresponding one-dimensional simplex. At any points, the connection 1-form is determined by continuity.
Hence, we assume that the fields $\omega_{\mu}$ smoothly depend on the points belonging to the geometric realization of each four-dimensional simplex. In this case, the following formula is valid up to $O\big{(}(\mathop{\rm d}\nolimits x)^{2}\big{)}$ inclusive
$\Omega_{{{\cal V}}_{m}{{\cal V}}_{i}}\Omega_{{{\cal V}}_{i}{{\cal V}}_{j}}\Omega_{{{\cal V}}_{j}{{\cal V}}_{m}}=\mathop{\rm exp}\nolimits\left[\frac{1}{2}\,\mathop{{\mathfrak{R}}}\nolimits_{\mu\nu}(x_{{{\cal V}}_{m}})\mathop{\rm d}\nolimits x^{\mu}_{\left({{\cal V}}_{m}{{\cal V}}_{i}\right)}\mathop{\rm d}\nolimits x^{\nu}_{\left({{\cal V}}_{m}{{\cal V}}_{j}\right)}\right],$ (31)
where
@@
In exact analogy with (26), let us write out the following relations for a tetrad field without explanations:
$\hat{e}_{\mu}\left({\cal W},{{\cal V}}_{m},{{\cal V}}_{i}\right)\mathop{\rm d}\nolimits x^{\mu}_{{{\cal V}}_{m}{{\cal V}}_{i}}=\hat{e}_{{{\cal V}}_{m}{{\cal V}}_{i}},\quad i=1,2,3,4.$ (33)
So 1-form
$e^{a}(x)=e^{a}_{\mu}(x)\mathop{\rm d}\nolimits x^{\mu}$ (34)
arises.
With the help of Equations given in this Section, we rewrite the lattice action (4) in the long-wavelength limit:
@@
In formulas (35)-(37) the additional subscript "$Lat"$ means that the corresponding values are defined on the lattice.
In the long-wave limit, there is a general covariance of the theory: all terms of the action (35)-(37) are invariant relative to general coordinate transformations.

### v9

@@ Abstract
Lattice regularization of the theory of gravity provides a new possibility for study of the Big Bang problem. In the 4D lattice theories of gravity, the existence of a high-temperature phase is proved, which is characterized by the folding of space into a point and zeroing of the energy-momentum tensor. The existence of a low-temperature phase in the long-wave limit is also shown, the geometric properties of which and its dynamics correspond to known concepts: the expansion of the Universe first follows an exponential law, and then smoothly transitions to a power-law regime.
@@ I Introduction
As is known, in the Minimal Standard Model (MSM) the number of fermionic degrees of freedom is multiple than the number of bosonic degrees of freedom (even taking into account two gravitons). It follows that the total energy of vacuum zero-point oscillations is negative. We assume that some properties of the MSM near the point of the Big Bang can be described using the discrete gravity model being studied. Here we are interested in the following question: can the continuation of solutions to Einstein’s continuum equations in the vicinity of the Big Bang point contain some information about the properties of the high-temperature phase in the lattice theory? To answer this question, we solve Einstein’s equations in the spirit of Friedmann’s model, but with a bare cosmological constant of positive sign. The positive cosmological constant compensates for the enormous negative energy of the vacuum. And since we consider the continuum theory as the long-wave limit of the discrete theory of gravity, then all the quantities in the equations (vacuum energy density $\varepsilon$, the bare cosmological constant $\Lambda_{0}$, etc.) are finite. In this case, Einstein’s equations and their solutions are mathematically correct. Moreover, the solutions obtained demonstrate the necessary properties common to cosmology: near the point of the Big Bang, an exponential expansion regime of the Universe takes place (inflation phase), but then the expansion switches to a power-law regime. It is also interesting that in the solution found, the vacuum energy modulus $|\varepsilon|$ decreases as we approach the Big Bang point. In other words, the energy-momentum tensor of matter in the continuum theory tends to zero when approaching the Big Bang point. In our opinion, this trend suggests that the Einstein continuum equation under consideration actually models a discrete theory of gravity in the long-wave limit. However, one must keep in mind that as the absolute value of the energy-momentum tensor decreases, the role of quantum fluctuations increases, and the classical Einstein equations become inapplicable (see Section IV).
@@
Section IV considers the solution of the Einstein equation within the Friedmann paradigm in the long-wave limit. This consideration makes sense, since the lattice theory transforms into the ordinary theory of gravity in the long-wave limit. It is shown that the solution of classical equations is possible only at some distance from the point of the Big Bang. The reason for this is that as one approaches the singularity, quantum fluctuations become significant.
@@ II.1 Lattice theory definition
This subsection uses the Euclidean signature. Denote by $\gamma^{a}$ $4\times 4$ Hermitian Dirac matrices, so that
@@
Consider the orientable abstract 4-dimensional simplicial complex $\mathop{{\mathfrak{K}}}\nolimits$. Suppose that any of its 4-simplexes belongs to such a finite (or infinite) sub-complex ${\mathop{{\mathfrak{K}}}\nolimits}^{\prime}\in\mathop{{\mathfrak{K}}}\nolimits$ which has a geometric realization in $\mathbb{R}^{4}$ topologically equivalent to a disk without cavities. The vertices are designated as $a_{{\cal V}}$, the indices ${{\cal V}}=1,2,\dots,\,{\mathop{{\mathfrak{N}}}\nolimits}^{(0)}\rightarrow\infty$ and ${{\cal W}}$ enumerate the vertices and 4-simplices, correspondingly. It is necessary to use the local enumeration of the vertices $a_{{\cal V}}$ attached to a given 4-simplex: the all five vertices of a 4-simplex with index ${{\cal W}}$ are enumerated as $a_{{{\cal V}}_{({{\cal W}})i}}$, $i=1,2,3,4,5$. The later notations with extra low index $({{\cal W}})$ indicate that the corresponding quantities belong to the 4-simplex with index ${{\cal W}}$. Note that a quantity belonging to the 4-simplex $s^{4}_{{\cal W}}$ can also belong to the neighboring 4-simplex $s^{4}_{{\cal W}^{\prime}}$ if these 4-simplices have common simplices of lower dimension. The Levi-Civita symbol with in pairs different indexes $\varepsilon_{{{\cal V}}_{({{\cal W}})1}{{\cal V}}_{({{\cal W}})2}{{\cal V}}_{({{\cal W}})3}{{\cal V}}_{({{\cal W}})4}{{\cal V}}_{({{\cal W}})5}}=\pm 1$ depending on whether the order of vertices $s^{4}_{{\cal W}}=a_{{{\cal V}}_{({{\cal W}})1}}a_{{{\cal V}}_{({{\cal W}})2}}a_{{{\cal V}}_{({{\cal W}})3}}a_{{{\cal V}}_{({{\cal W}})4}}a_{{{\cal V}}_{({{\cal W}})5}}$ defines the positive or negative orientation of 4-simplex $s^{4}_{{\cal W}}$. An element of the compact group $\mathop{\rm Spin}\nolimits(4)$ and an element of the Clifford algebra
$\Omega_{{{\cal V}}_{1}{{\cal V}}_{2}}=\Omega^{-1}_{{{\cal V}}_{2}{{\cal V}}_{1}}=\mathop{\rm exp}\nolimits\left(\omega_{{{\cal V}}_{1}{{\cal V}}_{2}}\right)=\mathop{\rm exp}\nolimits\left(\frac{1}{2}\sigma^{ab}\omega^{ab}_{{{\cal V}}_{1}{{\cal V}}_{2}}\right)\in\mathop{\rm Spin}\nolimits(4),$
@@
$\mathop{{\mathfrak{A}}}\nolimits_{g}=-\frac{1}{5!\cdot 2\cdot l_{P}^{\prime 2}}\sum_{{\cal W}}\sum_{\sigma}\varepsilon_{\sigma({{\cal V}}_{({{\cal W}})1})\sigma({{\cal V}}_{({{\cal W}})2})\sigma({{\cal V}}_{({{\cal W}})3})\sigma({{\cal V}}_{({{\cal W}})4})\sigma({{\cal V}}_{({{\cal W}})5})}$
@@
$\mathop{{\mathfrak{A}}}\nolimits_{\Lambda_{0}}=-\frac{1}{5!\cdot 12}\cdot\frac{\Lambda_{0}}{l_{P}^{\prime 2}}\varepsilon_{abcd}\sum_{{\cal W}}\sum_{\sigma}\varepsilon_{\sigma({{\cal V}}_{({{\cal W}})1})\sigma({{\cal V}}_{({{\cal W}})2})\sigma({{\cal V}}_{({{\cal W}})3})\sigma({{\cal V}}_{({{\cal W}})4})\sigma({{\cal V}}_{({{\cal W}})5})}$
@@ II.2 The long-wavelength limit
Let’s pass on to the long-wave limit, that is, to the limit of fields that slowly change when moving along the lattice. In this limit, the action (4) transforms into the well-known continuum Palatini action plus the contribution of the cosmological constant, and the Dirac field turns out to be minimally coupled to gravity. This transformation have meaning together with the transition to Minkowski signature. As a result the compact gauge group $\mathop{\rm Spin}\nolimits(4)$ transforms into the non-compact group $\mathop{\rm Spin}\nolimits(3,1)$. In this section, all lattice quantities for the Euclidean signature are primed. For the quantities related to the Minkowski signature, the previous designations are retained.
For the indicated transformation of the action, the following deformations of the integration contours in the integral (10) are necessary:
@@
The transition to the long-wave limit is possible for such field configurations that change quite slowly during transitions from simplex to simplex, that is, with small or significant movements along the lattice. This rule applies to any grating. In our theory, it is precisely at the stage of transition to the long-wave limit that the need to introduce local coordinates arises. Local coordinates are the markers of the lattice vertices. Consider some 4D subcomplex ${\mathop{{\mathfrak{K}}}\nolimits}^{\prime}\in\mathop{{\mathfrak{K}}}\nolimits$ with the trivial topology of a four-dimensional disk and geometric realization in $\mathbb{R}^{4}$. Thus, each vertex of the subcomplex acquires coordinates $x^{\mu}$, which are the coordinates of the image of the vertex in $\mathbb{R}^{4}$:
@@
At this stage, the coordinates are dimensionless, that is, real numbers. Let’s consider some simplex $s^{4}_{{\cal W}}\in{\mathop{{\mathfrak{K}}}\nolimits}^{\prime}$. Let us denote all five vertices of this 4-simplex as ${{\cal V}}_{i}$, $i=1,2,3,4$ and ${{\cal V}}_{m}\neq{{\cal V}}_{i}$. The properties of the geometric realization are such that four infinitesimal vectors
@@
are linearly independent.
In the works Vergeles (2006, 2015, 2017a, 2017b, 2021a) it is proven that there are 1-forms in $\mathbb{R}^{4}$ $\omega_{\mu}(x)$ and $\hat{e}_{\mu}(x)$ such that the equalities
$\omega_{\mu}\left(\frac{1}{2}\,(x_{{{\cal V}}_{m}}+x_{{{\cal V}}_{i}})\,\right)\mathop{\rm d}\nolimits x^{\mu}_{{{\cal V}}_{m}{{\cal V}}_{i}}=\omega_{{{\cal V}}_{m}{{\cal V}}_{i}},$ (26)
$\hat{e}_{\mu}\left(\frac{1}{2}\,(x_{{{\cal V}}_{m}}+x_{{{\cal V}}_{i}})\,\right)\mathop{\rm d}\nolimits x^{\mu}_{{{\cal V}}_{m}{{\cal V}}_{i}}=\hat{e}_{{{\cal V}}_{m}{{\cal V}}_{i}}.$ (27)
are true. It is further assumed that the 1-forms smoothly depend on points in $\mathbb{R}^{4}$ that are images of the vertices of the complex. In the long-wave limit, the elements $\Omega_{{{\cal V}}_{m}{{\cal V}}_{i}}$ are close to unity and, up to $O\big{(}(\mathop{\rm d}\nolimits x)^{2}\big{)}$ inclusive, the formula
$\Omega_{{{\cal V}}_{m}{{\cal V}}_{i}}\Omega_{{{\cal V}}_{i}{{\cal V}}_{j}}\Omega_{{{\cal V}}_{j}{{\cal V}}_{m}}=\mathop{\rm exp}\nolimits\left[\frac{1}{2}\,\mathop{{\mathfrak{R}}}\nolimits_{\mu\nu}(x_{{{\cal V}}_{m}})\mathop{\rm d}\nolimits x^{\mu}_{\left({{\cal V}}_{m}{{\cal V}}_{i}\right)}\mathop{\rm d}\nolimits x^{\nu}_{\left({{\cal V}}_{m}{{\cal V}}_{j}\right)}\right]$ (28)
holds, where
@@
Let’s write down the long-wave limit of the action (4):
@@
All other terms in such a transition will contain additional factors in the positive power $(l_{P}/\lambda)\longrightarrow 0$, and therefore they are omitted. Here $\lambda$ is the characteristic wavelength of the physical subsystem. This situation is typical when going to the long-wave limit in any lattice theory.
For clarity, let us point out the fact that on the lattice all variables and constants are dimensionless and of order one. In particular, the constant $l_{P}^{\prime}\sim 1$ in (5) and (9) is dimensionless, as are the differentials $\mathop{\rm d}\nolimits x^{\mu}_{{{\cal V}}_{m}{{\cal V}}_{i}}$ in (25). When moving to dimensional quantities, we assume
$\mathop{\rm d}\nolimits x^{\mu}_{{{\cal V}}_{m}{{\cal V}}_{i}}=\mathop{\rm d}\nolimits x^{\mu}/l_{P}\sim 1,$ (33)
where the differential $\mathop{\rm d}\nolimits x^{\mu}$ is measured in centimeters and $l_{P}\sim 10^{-32}\mbox{cm}$ (see (53)). From (33) it is clear that the step of the irregular lattice has a size of the order of $l_{P}$. And at the same time, all components of the action are dimensionless, but variables and constants acquire dimension. For example, in the term (32) the cosmological constant is $\Lambda_{0}\sim l_{P}^{-2}$.
The action (30)-(32) is a Hilbert-Einstein action, minimally related to the Dirac field and written in Palatini form. It is invariant under diffeomorphisms. This fact is not accidental, since in (24) the very method of introducing coordinates is such that already at this stage the independence of the action from the arbitrariness of introduced coordinates is visible. It is important that small terms in the long-wave limit, proportional to positive powers of $(l_{P}/\lambda)$, are also invariant under diffeomorphisms. Indeed, when expanding the action (4) in terms of the parameter ł$\l_{P}$, which does not depend on the coordinates, all the coefficients at the powers of this parameter will also be independent on the coordinates.